OpenStudy (anonymous):

The midpoint of CD is E(–1, 0). One endpoint is C (5, 2). What are the coordinates of the other endpoint

OpenStudy (gorv):

let one end point is (a,b) & other is (5,2) now mid point is -1=(a+5)/2 -2=a+5 -7=a 0=(b+2)/2 b+2=0 b=-2

OpenStudy (gorv):

@historylover

OpenStudy (anonymous):

so is that the final answer? im sorry im just really confused

OpenStudy (gorv):

tell me ...????

OpenStudy (gorv):

(-7,-2) is final ans

OpenStudy (gorv):

@historylover

OpenStudy (anonymous):

thank you!